OpenStudy (anonymous):

Let A be a 2x2 matrix whose eigenvalues are 2 and -3. What is the determinant of A^3 + 2A^2- 5A + 3i, where i is the identity matrix.

OpenStudy (anonymous):

this question is just too awesome isn't it?

OpenStudy (zarkon):

actually pretty easy...try diagonalization

OpenStudy (anonymous):

I don't see how that will work out....

OpenStudy (anonymous):

From the characteristic poly, I can find detA = 6 I also know det(A-2i) = det(A+3i) = 0

OpenStudy (zarkon):

you don't need that

OpenStudy (zarkon):

write \[A=PDP^{-1}=P\left[\begin{matrix}2 & 0\\ 0 & -3\end{matrix}\right]P^{-1}\]

OpenStudy (anonymous):

I got that as well, though I don't know what to do with it...

OpenStudy (zarkon):

recall that \[A^n=PD^nP^{-1}=P\left[\begin{matrix}2 & 0\\ 0 & -3\end{matrix}\right]^nP^{-1}\] \[=P\left[\begin{matrix}(2)^n & 0\\ 0 & (-3)^n\end{matrix}\right]P^{-1}\]

OpenStudy (zarkon):

\[ A^3 + 2A^2- 5A + 3I\] \[= (PDP^{-1})^3 + 2 (PDP^{-1})^2- 5 (PDP^{-1}) + 3 (PP^{-1})\] \[= P[D^3 + 2 D^2- 5 D + 3 ]P^{-1}\]

OpenStudy (zarkon):

forgot my I \[= P[D^3 + 2 D^2- 5 D + 3I ]P^{-1}\]

OpenStudy (anonymous):

so det[A^3 + 2A^2- 5A + 3i] = det(P) + det [D^3 + 2D^2 - 5D + 3i] + det(P^-1)?

OpenStudy (zarkon):

\[Det\left(P[D^3 + 2 D^2- 5 D + 3I ]P^{-1}\right)\] \[=Det(P)\cdot Det\left[D^3 + 2 D^2- 5 D + 3 I]\right]\cdot Det(P^{-1})\]

OpenStudy (anonymous):

o right lol, multiply not add, but also...

OpenStudy (anonymous):

the P and P^-1 cancel out leaving with det[D^3 + 2D^2 - 5D + 3i] which is very obvious to figure out at that point...

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

thx a lot, been pointed at the wrong direction for a long time now lol, thank goodness you are here :D